### v1

# SVG360: Multi-View SVG Generation with Geometric and Color Consistency from a Single SVG

Abstract
Scalable Vector Graphics (SVGs) are central to modern design workflows, offering scaling without distortion and precise editability. However, for single object SVGs, generating multi-view consistent SVGs from a single-view input remains underexplored. We present a three stage framework that produces multi-view SVGs with geometric and color consistency from a single SVG input. First, the rasterized input is lifted to a 3D representation and rendered under target camera poses, producing multi-view images of the object. Next, we extend the temporal memory mechanism of Segment Anything 2 (SAM2) to the spatial domain, constructing a spatial memory bank that establishes part level correspondences across neighboring views, yielding cleaner and more consistent vector paths and color assignments without retraining. Finally, during the raster to vector conversion, we perform path consolidation and structural optimization to reduce redundancy while preserving boundaries and semantics. The resulting SVGs exhibit strong geometric and color consistency across views, significantly reduce redundant paths, and retain fine structural details. This work bridges generative modeling and structured vector representation, providing a scalable route to single input, object level multi-view SVG generation and supporting applications such as asset creation and semantic vector editing.
1 Introduction
Editable vector graphics are central to modern design workflows, providing resolution independence, structural clarity, and precise editability. However, generating multi view consistent and fully editable vector graphics for a given object remains an open challenge. When designers attempt to present an existing vector illustration from a new viewpoint, they often need to manually redraw or adjust its structure to approximate the desired perspective. This manual process is both time consuming and prone to geometric and stylistic inconsistencies. Adobe Illustrator’s Turntable (Beta) [1] is one of the few publicly available end-to-end tools that generate multiple views from a single 2D vector artwork, yet it provides constrained angle control and discloses no technical specification.
The task of automatically generating multi-view consistent and fully editable SVGs, however, has not been extensively explored. We define this task as synthesizing a set of path level editable SVGs that preserve geometric and stylistic coherence across viewpoints. Solving this problem is particularly valuable for creative workflows such as turntable style visualization, multi-view logo and icon generation, and geometry consistent SVG asset editing. Bridging the gap between generative modeling and structured vector representation is therefore critical for enabling scalable and reliable design pipelines.
Recent advances in generative AI, particularly diffusion models[10, 29, 25, 30, 26, 46, 23], have enabled photorealistic and view-consistent image synthesis from a single input view, with representative approaches including a series of image-conditioned multi-view generation models [20, 33, 21, 38].
Yet, these outputs remain raster based and lack the structured semantics and fine grained editability inherent to vector graphics. Meanwhile, existing vectorization approaches such as DiffVG [18], DeepSVG [3], and Im2Vec [28] focus on single view reconstruction and fail to maintain geometric consistency across views. This gap motivates a core objective: enabling generative models to produce multi view coherent and structurally editable representations directly in the SVG domain.
To realize this objective, we propose a unified generative to vector framework that produces multi-view, geometrically consistent, fully editable SVGs from a single input. Our method consists of three stages. It first performs three dimensional aware multi-view generation to obtain geometry plausible novel views. Then it performs cross-view refinement through spatially aligned segmentation and appearance harmonization, enhancing the consistency of object contours and colors across views in the SVG style domain. Finally, leveraging the multi-view consistent segmentation results, we perform raster-to-vector conversion followed by path consolidation and color correction, yielding compact, coherent, and easily editable multi-view SVGs.
On multi-view SVG generation, our method produces noticeably cleaner and more stable results than existing vectorization tools. Compared with Adobe Illustrator’s Turntable, the path counts of our generated views stay much closer to the structure of the input SVG, reducing the path-count deviation from the input SVG by 26.5%. The variation in color usage between adjacent viewpoints is also reduced by 83.5%, indicating a far more consistent preservation of the input color design across rotations. In addition, our method lowers the average number of paths by 11.6%, resulting in lighter and more editable multi-view SVG sequences. The generated views exhibit coherent appearance transitions and show strong potential for creative design workflows, turntable-style visualization, and semantically aligned vector editing. Our contributions are summarized as follows:
Unified Generative to Vector Framework. We introduce the first end-to-end framework to directly synthesize multi view consistent and fully editable SVGs from a single input, bridging generative modeling with structured vector representation.
Spatially-Aware Cross-View Segmentation Refinement. We design a spatially-aware segmentation propagation mechanism that replaces temporal adjacency with geometry-guided neighborhood traversal, ensuring consistent part-level segmentation and enhancing SVG generation accuracy across multiple views.
Structure Preserving Vector Consolidation. We develop a vector domain consolidation strategy that merges redundant paths while maintaining semantic boundaries and structural coherence, yielding compact and semantically stable SVG assets suitable for real world design workflows.
The remainder of this paper is organized as follows: Section 2 reviews related work on multi-view SVG generation and the open gap; Section 3 presents our three stage framework; Section 4 presents experimental comparisons and ablation studies; Section 5 concludes and discusses limitations.
2 Related Work
2.1 Single image to multi view generation
Turning one reference view into a usable turntable sequence has become a key precursor to our goal of editable multi-view SVGs. Existing approaches fall into two major categories:
Diffusion-based multi-view image synthesis builds on high-fidelity 2D diffusion priors. These methods condition the denoising trajectory on target camera poses to synthesize each view directly in the image domain. Representative works include Zero-1-to-3 [20] and Zero123++ [33], where the former targets novel-view images rather than explicit 3D, and the latter improves cross-view consistency through better conditioning and training. Further, SyncDreamer [21] explicitly couples views during sampling via synchronized diffusion to align geometry and color, and SV3D [38] leverages video-latent diffusion to generate orbital view sequences from a single image. Despite strong appearance fidelity, residual independence across views often leads to geometric and color drift on thin structures and self occlusions.
3D reconstruction and generation from single images. These methods predict a 3D asset from one or a few views and render all novel viewpoints from the shared geometry, which naturally improves multi-view geometric consistency over 2D diffusion synthesis. Representative approaches include mesh generators such as GET3D, PolyGen, MeshDiffusion [8, 24, 22], single-image reconstruction methods like LRM, TripoSR, UniQue3D [11, 36, 40], and latent 3D models including Shape-E and Trellis [14, 41]. However, their outputs remain rasterized images without explicit vector paths or cross-view path correspondences, making them unsuitable for SVG-oriented workflows.
3D Gaussian Splatting (3DGS) as a shared representation. 3DGS [15] represents a scene with anisotropic Gaussians and renders by fast differentiable splatting, so all views reproject the same underlying 3D field. This often yields strong cross view geometric stability and efficient multi view rendering. Recent works combine diffusion priors with 3DGS for rapid 3D asset creation [35, 7], and single image variants enable ultra fast inference [34]. Although 3DGS provides efficient and geometrically consistent multi-view rendering, its outputs remain raster-based and do not explicitly expose vector path topology. Consequently, converting rendered contours into clean, closed, and editable paths requires an additional vectorization stage.
2.2 Raster to Vector Generation
Transforming multi-view consistent rasters into structured vector representations remains challenging because recovering continuous topology and geometry from discrete pixels is inherently ill-posed. Classical contour tracing methods such as Potrace and AutoTrace detect boundaries and fit Bézier curves or polygons. They perform well on icons or low-color imagery, but on smooth gradients and complex textures they often produce jagged, fragmented, or redundant paths.
Learning-based approaches model vector paths parametrically. DiffVG [18] introduces a differentiable renderer for gradient-based optimization, and DeepSVG [3] and Im2Vec [28] leverage Transformer or VAE designs to improve geometric continuity and structural controllability. LIVE [45] decomposes images into hierarchical layers to preserve global topology, yet can over-segment fine details. More recent methods such as VectorFusion [13] and StarVector [39] incorporate diffusion priors or region fusion to obtain smoother and more coherent paths.
Despite these advances, all vectorization methods still operate directly on pixel-level color transitions, which makes complex objects prone to fragmented or redundant paths. Introducing semantic segmentation before vectorization greatly simplifies the problem by isolating coherent part regions, enabling cleaner per-part conversion and more stable geometric structures. This highlights the need for a semantic layer, especially when extending vectorization to multi-view settings.
2.3 Segmentation and Cross View Consistency
While semantic segmentation simplifies single-view vectorization by isolating coherent part regions, multi-view SVG generation raises an additional requirement: achieving more consistent part assignments across viewpoints greatly improves cross-view coherence and visual stability. Yet conventional segmentation relies mainly on per-image pixel cues, making such consistency difficult; a spatially stable segmentation step is therefore needed before vectorization to provide more reliable part structures across views.
Early two-dimensional methods [9, 4, 42, 17] focus on per-image quality but do not enforce label correspondence across different views of the same object, causing boundaries to drift or merge inconsistently after projection. Foundation-scale models such as SAM, HQ-SAM, and SAM2 [16, 5, 27] provide accurate boundaries and promptable interaction, thus serving as strong building blocks for multi-view scenarios. However, their propagation mechanisms are primarily temporal rather than spatial, which limits their ability to maintain consistent part correspondences across viewpoints in static 3D scenes.
Inspired by video object segmentation, memory-based propagation [6, 43] maintains features across frames to keep masks consistent over time. By analogy, multi-view consistency can be formulated as a spatial sequence problem on the viewing sphere, where each camera view is treated as a pseudo-temporal step and neighboring views act as spatial references.
Existing work either stabilizes masks in raster space or improves single-view vectorization quality, but a bridge between cross-view semantic consistency and vector-level path correspondence remains underexplored. This gap motivates our approach, which first achieves part-aware, cross-view consistent segmentation and then leverages these semantic units to drive compact and editable vector representations across views.
3 Method
Figure 1: Our pipeline begins by converting the input SVG into a raster image, followed by rendering 3D consistent multi-view rasters using Trellis. A lightweight LoRA [12] tuned FLUX [2] model is then applied to harmonize their appearance. The refined rasters are processed by our Spatial SAM2 module, which replaces temporal adjacency in SAM2 with a spatial nearest-neighbor traversal on the viewing sphere. During segmentation propagation, the Spatial Memory Selector retrieves the geometrically most relevant memory entries. For the target view $t_{n-3}$, spatially adjacent views such as $t_{9}$, $t_{55}$, or $t_{n-6}$ may be closer than its temporal neighbor $t_{n-4}$, which helps maintain part-level consistency across viewpoints. The resulting masks are vectorized by VTracer [37] and further refined in the vector domain to produce compact, editable, and cross-view consistent multi-view SVGs.
As illustrated in Figure 1, our SVG360 framework begins with a 3D-aware rasterization stage (§3.1) that lifts a single-view SVG into geometrically plausible multi-view rasters. We then perform spatially aligned segmentation propagation3.2) to establish part-level correspondences across multi-view. Finally, the vector consolidation stage (§3.3) converts the segmented rasters into compact, fully editable multi-view SVGs with consistent geometry and color.
3.1 Multi-View Raster Generation and Harmonization
Multi-view raster generation. Generating accurate and geometrically consistent multi-view images from a single SVG input is challenging, as purely 2D generation approaches struggle to ensure structural coherence across views. To address this, we adopt a generative 3D-based approach to synthesize multi-view images.
Although several recent single-image 3D methods (e.g., Sparc3D [19] and Hi3DGen [44]) demonstrate strong geometric fidelity, they primarily focus on precise shape reconstruction and often do not provide fully textured, render ready outputs. Such geometry-only representations are unsuitable for our color consistent raster to SVG multi-view workflow. Considering open-source availability, generation efficiency, and the level of visual fidelity required by our task, we adopt Trellis [41] as the backbone 3D generator. Trellis offers a balanced trade-off between geometric accuracy and texture consistency, and remains efficient and practical for producing coherent multi-view rasters from a single SVG input.
Trellis represents assets in a unified structured latent space and decodes them into multiple 3D formats, including meshes and 3D Gaussians [41]. Its mesh decoder builds on FlexiCubes [32], predicting per-voxel signed-distance values and extracting a watertight surface from the zero-level isosurface. While both decoders yield high-fidelity results, converting the latent to a mesh and then rasterizing from many viewpoints introduces extra discretization and baking steps, which may accumulate appearance deviations relative to the input SVG style and incur nontrivial runtime overhead. For efficiency and fidelity, we directly use the 3D Gaussian decoder and render dense multi-view rasters via Gaussian splatting, which preserves the appearance statistics needed by our downstream segmentation and vectorization stages.
Raster harmonization.
While a 3D proxy preserves global geometric consistency across viewpoints, the reconstruction can still exhibit local errors, especially along edges and curved surfaces. View dependent illumination further introduces color variation that destabilizes raster-to-SVG conversion. We apply appearance harmonization using FLUX.1-dev [2] fine-tuned with a LoRA [12] trained on flat-design SVG data. The harmonized renders increase intra-region homogeneity and boundary contrast, suppress spurious shading edges, and reduce color variation across views. The resulting rasters are both view-consistent and faithful to SVG-style appearance, which improves downstream segmentation separability and stabilizes subsequent vector path consolidation.
3.2 Spatially Aligned Segmentation Propagation
After appearance harmonization, the multi-view rasters exhibit more uniform regions and sharper boundaries. To obtain cleaner and semantically stable vector paths in the subsequent vectorization stage, we need to maintain consistency of part-level identity across views. Relying on single view segmentation alone cannot enforce cross-view alignment of labels and boundaries, which weakens the robustness of path fitting and cross-view correspondence. Multi-view SVG generation inherently requires segmentation to be continuous in space. Accordingly, we spatialize the continuous segmentation paradigm of SAM2 by replacing temporal adjacency with spatial neighborhood on the viewing sphere and propagating masks along that sphere. Each target view is segmented under guidance from its nearest spatial neighbors, thereby maintaining part identity and boundary consistency at a global level.
Spatially Sequential Scheduling. We uniformly sample camera viewpoints on a unit sphere to obtain a globally distributed set of directions covering the entire view space. The sampling density and angular step are adjustable, allowing flexible control over the number of viewpoints and their spatial distribution according to task requirements. This configuration achieves globally balanced coverage even with a limited number of samples, forming a stable spatial adjacency structure that serves as the geometric foundation for subsequent cross-view propagation and visibility-aware reprojection. Although the sampled views have no intrinsic temporal order, a pseudo-sequential traversal helps maintain continuity during segmentation propagation. Starting from the front view aligned with the input SVG (denoted as $\theta_{0}$), each subsequent view is chosen as the unvisited one with the smallest angular distance from the current view, until all views are processed:
$\theta_{t+1}=\arg\min_{\theta\in\text{unvisited}}d(\theta_{t},\theta).$ (1)
This nearest-neighbor traversal minimizes geometric discontinuities between adjacent views, producing smoother and more stable mask propagation. To measure proximity between any two viewpoints, we define their angular distance as:
$d(\theta_{i},\theta_{j})=\mathrm{atan2}\big(\|u_{i}\times u_{j}\|,\,\mathrm{clip}(u_{i}\cdot u_{j},\,-1,\,1)\big),$ (2)
where $u(\psi,\phi)$ denotes the unit viewing direction for a camera defined by horizontal rotation (yaw, $\psi$) and vertical rotation (pitch, $\phi$):
$u(\psi,\phi)=[\cos\phi\cos\psi,\;\sin\phi,\;\cos\phi\sin\psi].$ (3)
This $\mathrm{atan2}$-based formulation is numerically more stable than $\arccos$, particularly when two directions are nearly parallel or opposite, and directly corresponds to the great-circle distance on the unit sphere. To further enhance traversal smoothness, we apply a lightweight two-segment swap (2-opt) optimization to remove local discontinuities and maintain continuous transitions between neighboring viewpoints.
Reference View Selection. After determining the traversal order, we construct a reference set for each target view to reuse information from geometrically adjacent and previously processed views. For each target view $\theta$, we select its $k$ nearest processed neighbors within an angular threshold $\tau$, ensuring that the guidance originates from spatially reliable local contexts. We empirically set $k=6$ and $\tau=75^{\circ}$ in all experiments. This localized reference selection, combined with the unified angular distance metric, restricts soft prompting and mask propagation to geometrically adjacent views, thereby preserving consistent part identities and boundary alignment across the entire viewing sphere.
Key Frame Initialization and Filtering.
Segmentation starts from the key view $\theta_{0}$, which corresponds to the input SVG. After obtaining initial masks from the automatic segmentation model, we apply a lightweight post-processing to reduce redundancy and improve stability for subsequent propagation. Small regions smaller than either 200 pixels or 0.05% of the image area (whichever is larger) are removed, a single morphological closing with an elliptical kernel sized to about 0.5% of the shorter image dimension is used to smooth boundaries, and overlapping masks are suppressed using non-maximum suppression with an IoU threshold of 0.5. The cleaned masks serve as prompt inputs to the spatially SAM2, providing stable guidance for the following multi-view segmentation process.
Residual Discovery.
Since the initial segmentation relies solely on the key view $\theta_{0}$, previously unseen parts that appear in other viewpoints may remain undetected. After Spatial-SAM2 completes segmentation across all sampled views, we introduce a residual discovery loop to automatically detect and recover these missing regions. For each processed view $\theta_{t}$, we compute the uncovered foreground ratio $r$ based on the union of all current masks. If $r$ exceeds 5% of the estimated foreground area, the same automatic segmentation procedure used for the key view is reactivated, but restricted to the uncovered regions. Newly discovered masks are then merged into the existing mask set and propagated to subsequent views through Spatial SAM2 to maintain cross-view consistency. The overall process is summarized in Algorithm 1.
Input: Set of views $\{\theta_{0},\dots,\theta_{N}\}$ with their current segmentation masks
Output: Updated mask set with newly discovered parts
$k\leftarrow 0$; // iteration counter
$T_{\max}\leftarrow 3$; // maximum number of passes
repeat
  for each view $\theta_{t}$ in spatial order do
     Compute uncovered foreground ratio $r$ using the union of all masks in $\theta_{t}$;
     if $r>5\%$ then
        Re-run automatic segmentation (same mode as key view), restricted to uncovered regions;
        Merge new masks into existing results;
        Propagate updates to neighboring views via Spatial-SAM2;
 $k\leftarrow k+1$;
until no new parts are found or $k\geq T_{\max}$;
Algorithm 1 Residual Discovery Loop
3.3 Vector Consolidation
Figure 2: Qualitative comparison. The figure summarizes representative issues observed in Adobe Turntable: (a) geometric inconsistencies across adjacent views; (b) cluttered structures arising from overlapping thin components; (c) merging of parts with similar colors; (d) missing regions in certain viewpoints; (e) color drift and gradual loss of small details across views. Our method produces multi-view SVGs with stable geometry, clear part separation, and consistent color appearance.
After segmentation, we crop each part into a transparent raster and feed it to VTracer [37], which converts pixel regions into path sets by color layering and spline fitting. Processing per part rather than per image reduces cross-part interference and yields a more compact path topology. This simplifies later consolidation and improves cross-view correspondence.
Direct vectorization often produces redundant small fills and micro strokes. We apply a light consolidation in the vector domain: sparsify colors using filled-area statistics to keep dominant colors and fold near-duplicates; clean only stroke-only micro paths to suppress artifacts from anti-aliasing; finally merge all parts within a view into a single SVG. These steps use fixed yet resolution-aware thresholds so the procedure is simple and efficient. This consolidation reduces the number of paths and colors, improves topological simplicity and editability, and yields more consistent boundaries across neighboring views.
To reduce small cross-view color drift, we extract a reference palette from the input SVG and map each source color to the nearest palette entry using the CIEDE2000 color-difference metric [31]. We add a near-black bias to stabilize dark tones: colors in low lightness and low chroma preferentially map to black, otherwise to the closest non-black reference color. This preserves semantic colors while suppressing banding due to shadows and reflections. In practice we preserve alpha values, resolve inherited attributes into explicit colors, and remove empty attributes so that rendering is consistent across editors. Together, color alignment unifies appearance and consolidation compresses structure, improving cross-view consistency and the editability of the final SVGs.
4 Experiments
4.1 Implementation Details
For the 3D generation stage, we employ the TRELLIS-image-large model to generate multi-view rasterizations from the input SVG. To harmonize the rendered appearance, lightweight LoRA adapter is trained to adapt FLUX.1-dev to SVG-style appearance, following standard PEFT settings (rank=32). The text encoder is frozen, and the adapter is trained for 1,000 steps. All experiments are conducted on a single NVIDIA A100-80GB PCIe GPU.
4.2 Evaluation Metrics
We use two metrics to capture cross-view stability. $\mathrm{RMSE}_{\text{path}}$ measures how much each view’s path count deviates from the input SVG. $\overline{\Delta N}_{\text{color,nbr}}$ measures the change in color count between adjacent views, lower indicates more stable color behavior.
4.3 Comparison
Baseline Adobe Illustrator Turntable (Beta) is a commercial, publicly available end-to-end feature that generates multi-view SVGs from a single 2D vector artwork. Its internal implementation, view grid, and technical specifications are not disclosed, and no programmatic interface is provided. We therefore use Turntable as the industrial baseline for our full pipeline comparison.
According to the official documentation, Turntable supports approximately $\pm 120^{\circ}$ horizontal rotation in yaw (sampled every $15^{\circ}$) and three pitch levels at $\{-45^{\circ},0^{\circ},+45^{\circ}\}$. For fair comparison, we adopt the same camera poses defined by the Turntable configuration in our experiments.
Because Turntable only allows manual per-view SVG export, we evaluate this baseline on 10 representative SVG examples from which complete view sequences can be consistently obtained.
Quantitative Comparison Table 1 reports the numerical statistics for Adobe Turntable and our method. Our approach produces fewer paths and colors per view, indicating cleaner vector topology and a more compact palette. For cross-view stability, our method achieves lower $\mathrm{RMSE}_{\text{path}}$ and a substantially lower $\overline{\Delta N}_{\text{color,nbr}}$, showing that the generated views remain closer to the original SVG structure and maintain more consistent color behavior across viewpoints. Together, these results confirm that our multi-view SVGs are structurally simpler and exhibit smoother appearance transitions.
Table 1: Comparison of multi-view SVG statistics between Adobe Turntable and our method.
Metric Adobe Turntable Ours
$\overline{N}_{\text{path}}$ 29.59 26.15
$\overline{N}_{\text{color}}$ 4.80 3.37
$\mathrm{RMSE}_{\text{path}}$  ( $\downarrow$ ) 16.37 12.03
$\overline{\Delta N}_{\text{color,nbr}}$  ( $\downarrow$ ) 2.12 0.35
Qualitative Comparison. Figure 2 summarizes the typical failure patterns we observe when applying Adobe Turntable to single-SVG multi-view generation. Across multiple objects and views, Adobe Turntable frequently exhibits geometric inconsistencies, structural confusion, unstable camera poses, view-dependent clipping, and noticeable color drift. In contrast, our method maintains coherent geometry, clear part separation, and consistent appearance across views.
In terms of geometry, object structures do not remain consistent across views. In example (a), components that should be rigid, such as the roof rack and the headlights—change noticeably between yaw $0^{\circ}$ and $90^{\circ}$, with the rack switching from two crossbars to a flat plate and the paired headlights collapsing into a single central light. Examples (b) and (c) further show that when the input contains multiple thin or closely spaced parts, the generated views often introduce overlaps, clutter, or unintended merging, making individual components difficult to distinguish.
Viewpoint-related inconsistencies are also apparent. In examples (a) and (e), the result at yaw $90^{\circ}$ deviates from a clean side view, exhibiting a slight tilt and geometric deformation. When the input SVG spans a large spatial extent, Turntable does not normalize its scale, causing portions of the object to be clipped outside the rendering canvas, as observed in both (a) and (e).
Color stability across views can also be unreliable. In example (e), the sesame seeds on the burger change brightness across viewpoints, and those on the far side gradually diminish as the camera rotates. In example (c), regions with similar color are merged into a single path, reducing structural separation.
In contrast, our method relies on explicit 3D representation, controlled camera poses, and consistent color and path processing. As a result, the generated multi-view SVGs maintain stable geometry, clear part separation, and coherent appearance across all viewpoints.
4.4 Ablation Study
Figure 3: Segmentation comparison of Spatial-SAM2 and the original SAM2 tracking mode at view 47. The bottom row shows the each six reference views used for memory support, with view 0 serving as the initial key frame. Results shown here correspond to the first iteration before any subsequent refinement.
Figure 3 illustrates the difference between the original Tracking SAM2(Video Segmentation) and our Spatial SAM2 when propagating masks across multi-view settings. The propagation mechanism of Tracking Mode SAM2 is fundamentally linear and one-directional: it progresses strictly along the view sequence and therefore relies on the temporally adjacent set of reference views (views 41–46). While such linear propagation is appropriate in video scenarios, it introduces a notable limitation in static multi-view renderings. If certain structures are not robustly tracked over several consecutive views, the resulting errors accumulate and become amplified along the propagation chain. Consequently, at views with large spatial offsets, the tracking mode frequently fails to recover local structures such as wheels, leading to partial or complete omission.
Our Spatial SAM2, in contrast, does not depend on single-direction propagation. Instead, it selects reference views based on true geometric proximity on the viewing sphere. Besides leveraging temporally adjacent views such as 44–46, it also incorporates views like 24–26, which may be far apart in the temporal order but remain spatially close to the target view. Since reference information is validated jointly across multiple spatial neighbors rather than flowing along a single temporal chain, Spatial SAM2 effectively suppresses error accumulation and enhances segmentation robustness. As a result, even at spatially distant viewpoints, structures such as wheels remain stable and consistently segmented across views.
Figure 4: Ablation study of four segmentation strategies: Spatial SAM2 (ours), SAM2 Tracking Mode, SAM2 Auto Mode, and a segmentation free VTracer baseline.
As shown in Figure 4, we further compare the impact of four segmentation strategies on the final SVG outputs. Spatial-SAM2 produces the most stable and coherent results. The tracking mode of SAM2 still exhibits occasional misses in certain viewpoints. The Auto mode (i.e., SAM2’s Segment Everything) tends to over-fragment continuous structures due to the lack of cross-view constraints, resulting in inconsistent topology across views. When segmentation is entirely omitted, many parts fail to survive the vectorization stage and cannot be reliably reconstructed, leaving only a few coarse contours. These findings confirm that enforcing spatially consistent segmentation is critical for producing high-quality multi-view SVGs.
5 Conclusion
We presented a unified generative-to-vector framework that produces geometrically consistent and fully editable multi-view SVGs from a single input illustration. By integrating 3D-aware multi-view generation, spatially aligned cross-view refinement, and vector-domain consolidation, our approach bridges raster-based synthesis with structured vector representations and enables practical multi-view SVG editing workflows.
Despite these improvements, several challenges remain. Part-level consistency is still constrained by the limitations of current segmentation models, particularly on objects with intricate topology or numerous fine-grained components. Our framework currently focuses on single-object inputs and does not yet extend to scene-level vector graphics requiring instance reasoning or world-coordinate camera transformations. Furthermore, vectorization remains most reliable for closed, well-formed paths, while open or stroke-only structures provide weaker cues for stable multi-view reconstruction.
These limitations suggest promising directions for future work, including stronger segmentation priors, more robust cross-view correspondence strategies, scene-level vector generation, and topology-aware handling of open-path SVGs.